6 years ago, rosa was 4 times as old as her daughter. 10 years from now, she will be only twice as old. how old are they now?
Answer by mananth(16946)

daughter age = x
6 years ago she was x-6
Rosa's age 6 years ago= 4(x-6)
..
10 years later daughter = x+10
Rosa will be 4*(x-6)+16
..
4*(x-6)+16=2*(x+10)
4x-24+16=2X+20
4x-2x=24-16+20
2x=28
divide by 2
2x/2 = 28/2
x= 14 daughter's age now.
10 years later Rosa's age = 4(x-6)+16
plug value of x
4(14-6)+16
=32+16
=48
.
Now Rosa's age will be 48-10
= 38.